# Enhancing Transport Barriers with Swimming Microorganisms in Chaotic Flows

### v1

@@ Abstract
We investigate the effects of bacterial activity on the mixing and transport properties of two-dimensional, time-periodic flows in experiments and in a simple model. We focus on the interactions between swimming E. coli and the flow Lagrangian Coherent Structure (LCS), which are computed from experimentally measured velocity fields. Experiments show that such interactions are non-trivial and can lead to transport barriers through which the tracer flux is significantly reduced. Using the Poincaré map, we show that these transport barriers coincide with the outermost members of elliptic LCSs known as Lagrangian vortex boundaries. Numerical simulations further show that elliptic LCSs can repel elongated swimmers and lead to swimmer depletion within Lagrangian coherent vortices. A simple mechanism shows that such depletion is due to the preferential alignment of elongated swimmers with the tangents of elliptic LCSs. Our results provide insights into understanding the transport of microorganisms in complex flows with dynamical topological features from a Lagrangian viewpoint.
@@ 1 Introduction
Many microorganisms live in environments characterized by currents (e.g., oceans, rivers, human intestines), and flow mediates many important microbial processes such as infection (Costerton et al., 1999; Yawata et al., 2016), nutrients uptake (Taylor & Stocker, 2012; Musielak et al., 2009), and reproduction (Riffell & Zimmer, 2007; Zimmer & Riffell, 2011). Flow exerts forces and torques on microorganisms that affect their motility dynamics and spatial distribution (Guasto et al., 2012; Wheeler et al., 2019). It also controls the transport of essential molecules including nutrients, oxygen, and signals of mates and predators (Kim et al., 2016). Microorganisms, in turn, can adapt their swimming behavior to these physical and chemical gradients (Stocker et al., 2008). Such interactions can lead to non-trivial phenomena such as rheotaxis (Hill et al., 2007; Marcos et al., 2012; Mathijssen et al., 2019), gyrotaxis (De Lillo et al., 2014; Borgnino et al., 2018), and chemotaxis (Locsei & Pedley, 2009; Stocker & Seymour, 2012), to name a few examples.
@@
Even simple shear flows, when coupled to cell motility and morphology, can give rise to complex transport and cell motility behavior (Rusconi et al., 2014; Durham et al., 2009; Ezhilan & Saintillan, 2015; Gustavsson et al., 2016). For instance, fluid shear can cause a torque that constantly rotates elongated cell body and can result in accumulation of motile bacteria and phytoplankton in high-shear-rate regions of the flow (Rusconi et al., 2014; Barry et al., 2015; Ezhilan & Saintillan, 2015). Near surfaces, shear flows can orient flagellated cells against the flow, causing bacteria and spermatozoa to swimming upstream (Tung et al., 2015; Zaferani et al., 2018; Mathijssen et al., 2019). Away from surfaces, shear gradients can trap bottom-heavy gyrotactic swimmers at certain depth of water column, causing the formation of intense cell assemblages called “thin layers” (Durham et al., 2009, 2013; Gustavsson et al., 2016). In unsteady and/or complex flows, transport of microorganisms shows intriguing phenomena such as aggregation, dispersion, and pattern formation (Torney & Neufeld, 2007; Khurana et al., 2011; Khurana & Ouellette, 2012; Zhan et al., 2014; Qin & Arratia, 2022), but are less understood. In time-periodic flows, simulations show that elongated swimmers can be trapped or repelled by elliptic islands depending on their shape and swimming speed (Torney & Neufeld, 2007); such trapping effects can lead to a reduction in long-term swimmer transport (Khurana et al., 2011). Recent experiments and simulations show that microswimmers can be trapped, repelled, or dispersed by vortices depending on the dimensionless path length and swimming speed (Qin & Arratia, 2022). In isotropic turbulence, simulations show that elongated swimmers, while remaining rather uniformly distributed, exhibit preferential alignment with instantaneous Eulerian fields such as local velocity (Borgnino et al., 2019), vorticity (Zhan et al., 2014) and velocity gradient (Pujara et al., 2018).
@@
Recently, it has been shown that Lagrangian Coherent Structure (LCS) can be a useful concept to understand the transport properties of swimming microorganisms in complex flows in both numerical simulations and experiments (Khurana & Ouellette, 2012; Dehkharghani et al., 2019; Ran et al., 2021; Si & Fang, 2021, 2022; Yoest et al., 2022). Simulations in chaotic flows show that elongated swimmers align with repelling LCSs of hyperbolic fixed points (Khurana & Ouellette, 2012), while later numerical studies show that elongated active particles have a much stronger alignment with attracting LCSs (Si & Fang, 2021, 2022), similar to passive elongated particles (Parsa et al., 2011). Experimental investigations that examine the interactions of swimming organisms and flow LCSs are few but show some intriguing phenomena. Experiments in model porous media show that bacteria align and accumulate near attracting LCSs and induce filamentous density patters (Dehkharghani et al., 2019). In time-periodic flows, experiments show that the accumulation of bacteria near attracting LCSs can attenuate stretching and hinder large-scale transport, although small-scale mixing is locally enhanced (Ran et al., 2021). Most, if not all, previous studies focus on attracting and repelling LCSs associated with the flow hyperbolic fixed points. That is not surprising since one expects large levels of strain near or around those fixed points. Less understood are swimmer interactions with elliptic LCSs, i.e., vortex-like flow dynamical features (Haller, 2015; Haller et al., 2016; Farazmand & Haller, 2016). That is the focus of this manuscript.
@@
Here, we experimentally investigate the effects of bacterial activity on the mixing and transport properties of a passive scalar in a time-periodic flow in experiments and simulations. We focus on the interaction of swimming bacteria (E. coli) with the elliptic LCSs of the flow. Results show that such interaction leads to transport barriers through which the fluxes of the passive tracer are significantly reduced. By constructing the Poincaré map from velocimetry data, we show that these transport barriers coincide with the outermost member of elliptic LCSs, or namely, Lagrangian vortex boundaries (LVBs). We further test these results in numerical simulations and find that elliptic LCSs repel elongated swimmers and lead to swimmer accumulation outside (or swimmer depletion inside) the Lagrangian vortices. A simple mechanism shows that the repulsion of swimmers is due to the preferential alignment of elongated swimmers with the tangents of elliptic LCSs. Overall, these results can be useful in understanding the transport of microorganisms in chaotic flows with elliptic dynamical features or nontrivial vortex structures.
[…]
@@ 2 Methods
[…]
@@ 3.1 Experimental Results
Figures 1(a) and 1(b) show sample snapshots of dye mixing in the buffer solution ($\phi_{b}=0\%$) and active suspension ($\phi_{b}=0.5\%$), respectively. Both snapshots are taken after $N=300$ periods of forcing, with dye initially confined on the right half of the images at $N=0$. The dye patterns reveal complex patterns and the underlying vortex structures, which are similar in the buffer and the active suspension. However, a main difference is the existence of undyed (void-like) regions near the center of vortex structures in the active suspension [Fig. 1(b)], while similar regions are dyed in the buffer case [Fig. 1(a)]. The undyed void-like regions near the vortex center appear primarily in the left half of the image that is initially not covered by dye. This suggests that the interplay of microbial activity and vortex structures leads to (enhanced) transport barriers through which the dye fails to penetrate or penetrates much more slowly. A video of the dye mixing processes in the buffer and the active suspension are shown in Movie 1.
To further understand the interaction between activity and vortex structures, we plot in Fig. 1(d) the flow Poincaré map obtained from PTV experiments. Lines in the map are stroboscopic particle trajectories that connect a particle’s initial position to its next position after a period ($T=5$ s). These trajectories are obtained by numerical integration of fluid particles in the experimentally measured velocity fields. The Poincaré map reveals several nested families of closed trajectories, whose behaviors mimic Kolmogorov–Arnold–Moser (KAM) tori in Hamiltonian dynamical systems (Ottino, 1989). These nested torus families define Lagrangian coherent vortices known as the elliptic LCSs, with each torus acting as a barrier that blocks tracer (dye) transport within the vortex (Katsanoulis et al., 2020; Haller et al., 2020; Aksamit & Haller, 2022). Unlike hyperbolic LCSs, elliptic LCSs are rotation-dominated regions that do not experience substantial (fluid) stretching. We note that Lagrangian coherent vortices are fundamentally different from the vortices defined by Eulerian fields such as vorticity. While we find a striking correspondence between the dye mixing patterns [Figs. 1(a) and 1(b)] and the Lagrangian coherent vortices in the Poincaré map [Fig. 1(d)], such similarity is absent for the Eulerian vorticity field in Fig. 1(c).
To quantitatively locate the elliptic LCSs, we calculate the trajectory rotation average (TRA) in the Poincaré map, as shown by the color code in Fig. 1(d). The TRA characterizes the average angular velocity of a particle trajectory and is defined as:
@@
Here, $\dot{\mathbf{x}}_{i}$ is a stroboscopic velocity related to the rate of change of a particle’s position in the Poincaré map, and $N$ is the number of periods over which the quantity is calculated. Although the TRA is originally defined for actual particle trajectories (Haller et al., 2021), we use it here for the stroboscopic trajectories in the extended phase space of the Poincaré map. We interpolate the values of the TRA onto a uniform spatial grid to obtain the TRA field (see Movie 2). In two-dimensional flows, elliptic LCSs can be located from the closed convex contours of the TRA field (Haller et al., 2021; Aksamit & Haller, 2022). Since elliptic LCSs are nested families, we identify the outermost convex TRA contours as the boundaries of Lagrangian coherent vortices, and the centroid of the innermost convex TRA contours as elliptic fixed points. For concision, we refer to the outermost members of elliptic LCSs as Lagrangian vortex boundaries (LVBs).
Figure 2: Enlarged photographs of the dye concentration field for (a) buffer and (b) active suspension, in the dashed square regions shown in Fig. 1. Green contours are Lagrangian vortex boundaries (LVBs) identified from the TRA field and red dots are elliptic fixed points. (c) Dye flux as a function of time $N$, for the vortex with the label “4”. Inset: average dye flux as a function of vortex Reynolds number $Re_{\Gamma}$, for the buffer (blue) and the active suspension (red). Different markers represent different vortices in the flow field. Error bars are standard deviations.
Figures 2(a) and 2(b) show Lagrangian vortex boundaries (LVBs) (green contours) and elliptic fixed point (red dots) identified from the TRA field, superimposed on the dye mixing patterns in the buffer and the active suspension, respectively. Although LVBs in chaotic flows are known to correspond to transport barriers (Katsanoulis et al., 2020; Haller et al., 2020; Aksamit & Haller, 2022), the LVBs in the buffer case [Fig. 2(a)] contain dye inside. On the other hand, the LVBs in the active case [Fig. 2(b)] enclose void-like undyed regions around the elliptic fixed points. This suggests that bacterial activity enhances the strength of LVBs as transport barriers in chaotic flows. We also notice some differences in the size and shape of LVBs for the buffer and active cases. Since the LVBs are calculated from experimentally measured velocity fields, this indicates that microbial activity also modifies the underlying velocity field.
To quantify the strength of the transport barriers, or the reduction in dye transport, we calculate dye flux into the Lagrangian vortices using the following mass balance equation:
@@
Here, $S(t)$ and $B(t)$ denotes the enclosed area and the contour of the LVBs, $C$ is the dye concentration field, $D$ is the dye diffusivity, $\mathbf{n}$ is the normal vector of LVBs, $\mathbf{v}$ is fluid velocity, and $\mathbf{v}_{B}$ is the velocity induced by the motion and deformation of LVBs. The first and second terms on the right hand side of Eq. 2 are the contributions to the dye flux by diffusion and convection, respectively. The left hand side of Eq. 2 gives a simple way to calculate the dye flux as the rate of change of the total dye enclosed by the LVBs. Figure 2(c) shows the dye flux as a function of time, for vortex with label “4” in the buffer and the active suspension. We find a significant reduction of dye flux into the LVBs with bacterial activity, especially at earlier times ($N<150$). At later times ($N>200$), the dye flux in the active case is larger than that in the buffer, as the result of a larger diffusive flux due to large dye concentration gradients.
We now explore the generality of the reduction of dye flux in active suspensions. We start by calculating the circulation of a Lagrangian vortex, defined as:
$\Gamma(t)=\oint\limits_{B(t)}\mathbf{v}\cdot\,\mathrm{d}\mathbf{l}=\int\limits_{S(t)}\omega_{z}\,\mathrm{d}A.$ (3)
Here, $\omega_{z}$ denotes the vorticity component in the out-of-plane direction. We then define the the vortex Reynolds number as: $Re_{\Gamma}=\left|\overline{\Gamma}\right|/\nu$, where $\overline{\Gamma}$ denotes the time-averaged circulation over a period, and the absolute value is included since $\overline{\Gamma}$ can be positive or negative depending on the sign of the vorticity. We calculate $Re_{\Gamma}$ for each Lagrangian vortex in Figs. 2(a) and 2(b); the labels correspond to the magnitude of their $Re_{\Gamma}$. Inset of Fig. 2(c) shows the averaged dye flux as a function of the vortex Reynolds number $Re_{\Gamma}$, for the buffer and active cases. We find that in both cases the average dye fluxes scale linearly with $Re_{\Gamma}$. This is because both the average dye flux and $Re_{\Gamma}$ are linearly proportional to the area of the Lagrangian vortex. More importantly, however, we find that the slope of the scaling is larger in the buffer case compared to the active case. This result shows quantitatively that bacterial activity reduces the dye flux into the Lagrangian vortices and therefore enhances the strength of transport barriers.

### v2

@@ Abstract
We investigate the effects of bacterial activity on the mixing and transport properties of a passive scalar in time-periodic flows in experiments and in a simple model. We focus on the interactions between swimming E. coli and the Lagrangian Coherent Structures (LCSs) of the flow, which are computed from experimentally measured velocity fields. Experiments show that such interactions are non-trivial and can lead to transport barriers through which the scalar flux is significantly reduced. Using the Poincaré map, we show that these transport barriers coincide with the outermost members of elliptic LCSs known as Lagrangian vortex boundaries. Numerical simulations further show that elliptic LCSs can repel elongated swimmers and lead to swimmer depletion within Lagrangian coherent vortices. A simple mechanism shows that such depletion is due to the preferential alignment of elongated swimmers with the tangents of elliptic LCSs. Our results provide insights into understanding the transport of microorganisms in complex flows with dynamical topological features from a Lagrangian viewpoint.
@@ 1 Introduction
Many microorganisms live in environments characterized by currents (e.g., oceans, rivers, human intestines), and flow mediates many important microbial processes such as infection (Costerton et al., 1999; Yawata et al., 2016), uptake of nutrients (Taylor & Stocker, 2012; Musielak et al., 2009), and reproduction (Riffell & Zimmer, 2007; Zimmer & Riffell, 2011). Flow exerts forces and torques on microorganisms that affect their motility dynamics and spatial distribution (Guasto et al., 2012; Wheeler et al., 2019). It also controls the transport of essential molecules including nutrients, oxygen, and signals of mates and predators (Kim et al., 2016). Microorganisms, in turn, can adapt their swimming behavior to these physical and chemical gradients (Stocker et al., 2008). Such interactions can lead to non-trivial phenomena such as rheotaxis (Hill et al., 2007; Marcos et al., 2012; Mathijssen et al., 2019), gyrotaxis (De Lillo et al., 2014; Borgnino et al., 2018), and chemotaxis (Locsei & Pedley, 2009; Stocker & Seymour, 2012), to name a few examples.
@@
Even simple shear flows, when coupled to cell motility and morphology, can give rise to complex transport and cell motility behavior (Rusconi et al., 2014; Durham et al., 2009; Ezhilan & Saintillan, 2015; Gustavsson et al., 2016). For instance, fluid shear can cause a torque that can rotate elongated cell body and can result in accumulation of motile bacteria and phytoplankton in high-shear-rate regions of the flow (Rusconi et al., 2014; Barry et al., 2015; Ezhilan & Saintillan, 2015). Near surfaces, shear flows can orient flagellated cells against the flow, causing bacteria and spermatozoa to swim upstream (Tung et al., 2015; Zaferani et al., 2018; Mathijssen et al., 2019). Away from surfaces, shear gradients can trap bottom-heavy gyrotactic swimmers at certain depth of water column, causing the formation of intense cell assemblages called “thin layers” (Durham et al., 2009, 2013; Gustavsson et al., 2016). In unsteady and/or complex flows, transport of microorganisms shows intriguing phenomena such as aggregation, dispersion, and pattern formation (Torney & Neufeld, 2007; Khurana et al., 2011; Khurana & Ouellette, 2012; Zhan et al., 2014; Qin & Arratia, 2022), but are less understood. In time-periodic flows, simulations show that elongated swimmers can be trapped or repelled by elliptic islands depending on their shape and swimming speed (Torney & Neufeld, 2007); such trapping effects can lead to a reduction in long-term swimmer transport (Khurana et al., 2011). Recent experiments and simulations show that microswimmers can be trapped, repelled, or dispersed by vortices depending on the dimensionless path-length and swimming speed (Qin & Arratia, 2022). In isotropic turbulence, simulations show that elongated swimmers, while remaining rather uniformly distributed, exhibit preferential alignment with instantaneous Eulerian fields such as local velocity (Borgnino et al., 2019), vorticity (Zhan et al., 2014) and velocity gradient (Pujara et al., 2018).
@@
Recently, it has been shown that Lagrangian Coherent Structure (LCS) can be a useful concept to understand the transport properties of swimming microorganisms in complex flows in both numerical simulations and experiments (Khurana & Ouellette, 2012; Dehkharghani et al., 2019; Ran et al., 2021; Si & Fang, 2021, 2022; Yoest et al., 2022). Simulations in chaotic flows show that elongated swimmers align with repelling LCSs of hyperbolic fixed points (Khurana & Ouellette, 2012), while later numerical studies show that elongated active particles have a much stronger alignment with attracting LCSs (Si & Fang, 2021, 2022), similar to passive elongated particles (Parsa et al., 2011). Experimental investigations that examine the interactions of swimming organisms and flow LCSs are few but show some intriguing phenomena. Experiments in model porous media show that bacteria align and accumulate near attracting LCSs and induce filamentous density patterns (Dehkharghani et al., 2019). In time-periodic flows, experiments show that the accumulation of bacteria near the attracting LCSs can attenuate stretching and hinder large-scale transport, although small-scale mixing is locally enhanced (Ran et al., 2021). Most, if not all, previous studies focus on attracting and repelling LCSs associated with the flow hyperbolic fixed points. That is not surprising since one expects large levels of strain near or around those fixed points. Less understood are swimmer interactions with elliptic LCSs, i.e., vortex-like flow dynamical features (Haller, 2015; Haller et al., 2016; Farazmand & Haller, 2016). That is the focus of this manuscript.
@@
Here, we experimentally investigate the effects of bacterial activity on the mixing and transport properties of a passive scalar in a time-periodic flow in experiments and simulations. We focus on the interaction of swimming bacteria (Escherichia coli) with the elliptic LCSs of the flow. Results show that such interaction leads to transport barriers through which the fluxes of the passive tracer are significantly reduced. By constructing the Poincaré map from velocimetry data, we show that these transport barriers coincide with the outermost member of elliptic LCSs, or namely, Lagrangian vortex boundaries (LVBs). We further test these results in numerical simulations and find that elliptic LCSs repel elongated swimmers and lead to swimmer accumulation outside (or swimmer depletion inside) the Lagrangian vortices. A simple mechanism shows that the repulsion of swimmers is due to the preferential alignment of elongated swimmers with the tangents of elliptic LCSs. Overall, these results can be useful in understanding the transport of microorganisms in chaotic flows with elliptic dynamical features or nontrivial vortex structures.
Figure 1: Photographs of the dye concentration field for chaotic mixing in (a) buffer solution ($\phi_{b}=0$) and (b) active suspension ($\phi_{b}=0.5\%$). Images are taken at $N=300$ periods after the start of the experiments. The imaged region is $60\times 60$ mm2; the scale bar represents 6 mm. The Reynolds number and path-length of the flow are $Re=7.2$ and $p=1.0$, respectively. (c) Vorticity (color code) and velocity (arrows) fields of the flow, corresponding the the first peak of a time period. (d) Poincaré map of the flow in the same imaged region as the dye field. The map is colored by the trajectory rotation average (TRA) calculated over 20 time periods.
@@ 2 Methods
Experiments are performed in the flow cell setup where a 2-mm thin conductive fluid layer is placed above an array of permanent magnets arranged in a disordered pattern (Voth et al., 2002; Ran et al., 2021, 2022). As a sinusoidal forcing (electrical current, 0.2 Hz frequency) is imposed on the fluid layer, the magnetic field induces a Lorentz force and creates spatially disordered vortex patterns [see flow field in Fig. 1(c)]. The resulting flow is characterized by two parameters: the Reynolds number and the path-length. The Reynolds number is defined as $Re=U\!L/\nu$, where $U=1.2$ mm/s is the average flow speed, $L=6.0$ mm is the characteristic length scale determined by the average spacing of the permanent magnets, and $\nu=1.0$ mm2/s is the fluid kinematic viscosity (water-like). The path-length is the normalized mean displacement of a typical fluid parcel in one forcing period, defined as $p=UT/L$, where $T=5$ s is the forcing period. Here, the Reynolds number and the path-length are $Re\approx 7.2$, $p\approx 1.0$; these conditions are known to lead to chaotic advection in this system (Ran et al., 2021; Voth et al., 2003).
Two main types of experiments are performed: dye mixing and particle tracking velocimetry (PTV). Dye mixing experiments are performed by labelling half of the fluid layer with a passive dye tracer ($6.25\times 10^{-5}~{}$ sodium fluorescein). Initially, the labeled and unlabeled portions are separated by a physical barrier. As the flow begins, the barrier is lifted, and the labeled fluid progressively penetrates the unlabeled portion with time. The dye concentration field is recorded by a CMOS camera (Flare 4M180) at 5 frames/s with a resolution of $2000\times 2000$ pixels. Particle tracking experiments are performed by seeding the fluid with 100-m large fluorescent polystyrene particles; the Stokes number of these fluorescent particles is $\mathcal{O}(10^{-4})$, indicating good tracer fidelity. Particle positions are recorded by the CMOS camera at 30 frames/s with a resolution of $1200\times 1200$ pixels. Particle trajectories are obtained using an in-house tracking algorithm (Crocker & Grier, 1996); these trajectories are then used to obtain the velocity fields from 6th-order polynomial fitting. Because the flow is time-periodic, we can combine particle positions at a given phase (relative to the forcing) to obtain up to 80,000 precise particle positions at each phase, which yields high spatial resolution (0.002 of the field of view), excellent temporal resolution (0.007 of a flow period), and velocities accurate to a few percent. Two different velocity maps are obtained from separate experiments, one in the presence and the other in the absence of bacteria.
Active suspensions are prepared by adding a strain (wild-type K12 MG1655) of Escherichia coli to an aqueous buffer solution of 2% KCl and 1% NaCl by weight; the swimming motility of E. coli does not seem to be affected by the salts (Ran et al., 2021). The strain of E. coli has a swimming speed of 10–20 m/s and a rod-shaped body of on average 2 m long and 0.5 m in diameter. We note that the Péclet numbers of the bacteria and the dye are both considerably large, with $Pe_{b}=U\!L/D_{b}\sim\mathcal{O}(10^{6})$, and $Pe_{d}=U\!L/D_{d}\sim\mathcal{O}(10^{4})$, where $D_{b}$ and $D_{d}$ are the effective diffusivities of the swimming E. coli and the dye, respectively. Bacteria are cultured in Luria–Bertani (LB Lennox, Sigma-Aldrich) liquid media at 37C overnight for 12 to 14 hours to attain a stationary phase of a number density of approximately $10^{9}$ cells/mL. The stationary-phase culture is centrifuged at 5,000 revolutions per minute for 3.5 minutes and resuspended into the buffer solution to attain a number density of $1.25\times 10^{10}$ cells/mL or a bacterial volume fraction of $\phi_{b}=0.5\%$. This bacterial volume fraction ($\phi_{b}=0.5\%$) is considered dilute (Kasyap et al., 2014; Ran et al., 2021), and large scale collective behavior/motion is not expected.
@@ 3.1 Experimental Results
Figures 1(a) and 1(b) show sample snapshots of dye mixing in the buffer solution ($\phi_{b}=0\%$) and active suspension ($\phi_{b}=0.5\%$), respectively. Both snapshots are taken after $N=300$ periods of forcing, with dye initially confined on the right half of the images at $N=0$. The concentration field reveals complex patterns and underlying vortex structures, which are similar in the buffer and the active suspension. However, a main difference is the existence of regions devoid of dye near the center of the vortex structures in the active suspension [Fig. 1(b)], while similar regions are dyed in the buffer case [Fig. 1(a)]. The regions devoid of dye near the center of the vortex appear primarily in the left half of the image that is initially not covered with dye. This suggests that the interplay of microbial activity and vortex structures leads to (enhanced) transport barriers through which the dye fails to penetrate or penetrates much more slowly. A video of the dye mixing processes in the buffer and the active suspension are shown in Movie 1.
To further understand the interaction between activity and vortex structures, we plot in Fig. 1(d) the flow Poincaré map obtained from PTV experiments. Lines in the map are stroboscopic particle trajectories that connect a particle’s initial position to its next position after a period ($T=5~{}$s). These trajectories are obtained by numerical integration of fluid particles in the experimentally measured velocity fields. The Poincaré map reveals several nested families of closed trajectories, whose behaviors mimic Kolmogorov–Arnold–Moser (KAM) tori in Hamiltonian dynamical systems (Ottino, 1989). These nested torus families define Lagrangian coherent vortices known as elliptic Lagrangian coherent structures (elliptic LCSs), with each torus acting as a barrier that blocks tracer (dye) transport within the vortex (Katsanoulis et al., 2020; Haller et al., 2020; Aksamit & Haller, 2022). Unlike hyperbolic LCSs, elliptic LCSs are rotation-dominated regions that do not experience substantial (fluid) stretching. We note that Lagrangian coherent vortices are fundamentally different from the vortices defined by Eulerian fields such as vorticity. While we find a striking correspondence between the dye mixing patterns [Figs. 1(a) and 1(b)] and the Lagrangian coherent vortices on the Poincaré map [Fig. 1(d)], such similarity is absent for the Eulerian vorticity field in Fig. 1(c).
To quantitatively locate the elliptic LCSs, we calculate the trajectory rotation average (TRA) on the Poincaré map, as shown by the color code in Fig. 1(d). The TRA characterizes the average angular velocity of a particle trajectory and is defined as:
@@
Here, $\dot{\mathbf{x}}_{i}$ is a stroboscopic velocity related to the rate of change of the position of a particle on the Poincaré map and $N$ is the number of periods over which the quantity is calculated. Although the TRA was originally defined for actual particle trajectories (Haller et al., 2021), we used it here for the stroboscopic trajectories in the extended phase space of the Poincaré map. We interpolate the values of the TRA onto a uniform spatial grid to obtain the TRA field (see Movie 2). In two-dimensional flows, elliptic LCSs can be located from the closed convex contours of the TRA field (Haller et al., 2021; Aksamit & Haller, 2022). Since elliptic LCSs are nested families, we identify the outermost convex TRA contours as the boundaries of Lagrangian coherent vortices, and the centroid of the innermost convex TRA contours as elliptic fixed points. For concision, we refer to the outermost members of elliptic LCSs as Lagrangian vortex boundaries (LVBs).
Figure 2: Enlarged photographs of the dye concentration field for (a) buffer and (b) active suspension, in the dashed square regions shown in Fig. 1. Green contours are Lagrangian vortex boundaries (LVBs) identified from the TRA field and red dots are elliptic fixed points. (c) Dye flow rate, $\mathbf{J}$, as a function of time $N$, for the vortex with the label “4”. Inset: time-averaged dye flow rate, $\overline{\mathbf{J}}$, as a function of the vortex Reynolds number $Re_{\Gamma}$, for the buffer (blue) and the active suspension (red). Different markers represent different vortices in the flow field. Error bars are standard deviations.
Figures 2(a) and 2(b) show the LVBs (green contours) and elliptic fixed point (red dots) identified from the TRA field, superimposed on dye mixing patterns in the buffer and the active suspension, respectively. Although LVBs in chaotic flows are known to be transport barriers for convection (Katsanoulis et al., 2020; Haller et al., 2020; Aksamit & Haller, 2022), the LVBs in the buffer case [Fig. 2(a)] contain dye inside due to diffusion. On the other hand, the LVBs in the active case [Fig. 2(b)] enclose regions devoid of dye near the elliptic fixed points. This suggests that bacterial activity enhances the strength of LVBs as transport barriers in chaotic flows. We also notice some differences in the size and shape of the LVBs for the buffer and active cases. Since the LVBs are calculated from experimentally measured velocity fields, this indicates that microbial activity also modifies the underlying velocity field.
To quantify the strength of the transport barriers, or the reduction in dye transport, we calculate the dye flow rate into the Lagrangian vortices using the following mass balance equation:
@@
Here, $S(t)$ and $B(t)$ denotes the enclosed area and the contour of the LVBs, $C$ is the dimensionless dye concentration field (normalized by the maximum fluorescence intensity), $D$ is the dye diffusivity, $\mathbf{n}$ is the normal vector of LVBs, $\mathbf{v}$ is fluid velocity, and $\mathbf{v}_{B}$ is the velocity induced by the motion and deformation of LVBs. The first and second terms on the right-hand side of Eq. 2 are the surface integrals of the dye fluxes through the LVBs due to diffusion and convection, respectively. The left-hand side of Eq. 2 gives a simple way to calculate dye flow rate as the rate of change of the total dye enclosed by the LVBs. Figure 2(c) shows the dye flow rate, $\mathbf{J}$, as a function of time, for vortex with label “4” in the buffer and the active suspension. We find a significant reduction of dye flow rate into the LVBs with bacterial activity, especially at earlier times ($N<150$). At later times ($N>200$), the dye flow rate in the active case is larger than that in the buffer, as a result of a larger diffusive flux due to large dye concentration gradients.
We now explore the generality of the reduction of dye flow rate in active suspensions. We start by calculating the circulation of a Lagrangian vortex, defined as
$\Gamma(t)=\oint\limits_{B(t)}\mathbf{v}\cdot\,\mathrm{d}\mathbf{l}=\int\limits_{S(t)}\omega_{z}\,\mathrm{d}A,$ (3)
where $\omega_{z}$ denotes the vorticity component in the out-of-plane direction. We then define the the vortex Reynolds number as: $Re_{\Gamma}=\left|\overline{\Gamma}\right|/\nu$, where $\overline{\Gamma}$ denotes the time-averaged circulation over a period, and the absolute value is included since $\overline{\Gamma}$ can be positive or negative depending on the sign of the vorticity. We calculate $Re_{\Gamma}$ for each Lagrangian vortex in Figs. 2(a) and 2(b); the labels correspond to the magnitude of their $Re_{\Gamma}$. Inset of Fig. 2(c) shows the time-averaged dye flow rate, $\overline{\mathbf{J}}$, as a function of the vortex Reynolds number $Re_{\Gamma}$, for the buffer and active cases. We find that in both cases the average dye flow rates scale linearly with $Re_{\Gamma}$. This is because both $\overline{\mathbf{J}}$ and $Re_{\Gamma}$ are linearly proportional to the area of the Lagrangian vortex. More importantly, however, we find that the slope of the scaling is larger in the buffer case compared to the active case. This result shows that the presence of bacteria enhances the barriers for scalar transport into the flow Lagrangian vortices.